# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

### v1

@@ C.1 Identical contacts
General formulas for the conductance and the noise intensity are obtained by averaging the expressions  (8), (9), (10) and (11) in the main text over the energy, using expressions for the amplitude (31). It is convenient to write the resulting formulas, using the parameters $\lambda,$ $R_{\theta}$ and $R_{\phi}$ instead of the parameters $t,$ $\theta$ and $\phi$, according to the following definitions:
@@
${\mathcal{T}}^{\text{HI}}=C\sum_{m,n=0}^{1}A_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,,$ (55)
${\mathcal{T}}_{2}^{\text{HI}}=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,,$ (56)
@@
×$\times(R_{\theta}(13-9R_{\theta}(1-R_{\phi})-17R_{\phi})+8R_{\phi})/2\,,$
$\tilde{A}_{(2,2)}=$ $-R_{\theta}(R_{\theta}(9-7R_{\theta}(1-R_{\phi})-11R_{\phi})+4R_{\phi})\,,$
@@ C.2 Non-identical contacts
$t_{\text{R,L}}$ $=te^{\pm\beta}\,,\quad|\beta|\leqslant\lambda$ (59)
$t_{\text{R}}t_{\text{L}}$ $=e^{-2\lambda}\,,$
$t_{\text{R}}/t_{\text{L}}$ $=e^{2\beta}\,.$
In this case we may still use the expressions (55), (38), but the coefficients should be redefined, $A_{(m,n)}\to A_{(m,n)}^{\text{u}}$, $\tilde{A}_{(m,n)}\to\tilde{A}_{(m,n)}^{\text{u}}$, as follows
$C^{\text{u}}=$ $C\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)\,,$ (60)
$A_{(0,0)}^{\text{u}}=$ $A_{(0,0)}\,,$
$A_{(0,1)}^{\text{u}}=$ $0\,,$
$A_{(1,0)}^{\text{u}}=$ $A_{(1,0)}+R_{\theta}\sinh^{2}\beta\,,$
$A_{(1,1)}^{\text{u}}=$ $A_{(1,1)}+2R_{\theta}\sinh^{2}\beta\,,$
$\tilde{C}^{\text{u}}=$ $\tilde{C}\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)^{2}\,,$
$\tilde{A}_{(0,0)}^{\text{u}}=$ $\tilde{A}_{(0,0)}+2R_{\theta}(2-R_{\theta})\sinh^{2}\beta\,,$
$\tilde{A}_{(0,1)}^{\text{u}}=$ $\tilde{A}_{(0,1)}\,,$
$\tilde{A}_{(1,0)}^{\text{u}}=$ $\tilde{A}_{(1,0)}+R_{\theta}(12-R_{\theta}(17-9R_{\theta})$
@@
$\tilde{A}_{(1,1)}^{\text{u}}=$ $\tilde{A}_{(1,1)}+R_{\theta}^{2}\sinh^{2}(2\beta)\,,$
$\tilde{A}_{(2,0)}^{\text{u}}=$ $\tilde{A}_{(2,0)}+R_{\theta}(R_{\theta}(1-R_{\phi})+R_{\phi})$
×$\times(R_{\theta}(13-7R_{\theta}(1-R_{\phi})-11R_{\phi})$
@@
$\tilde{A}_{(2,1)}^{\text{u}}=$ $\tilde{A}_{(2,1)}-4R_{\theta}(2-R_{\theta})(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}\beta$
$+10R_{\theta}^{2}(1-R_{\theta})(1-R_{\phi})\sinh^{4}\beta\,,$
$\tilde{A}_{(2,2)}^{\text{u}}=$ $\tilde{A}_{(2,2)}-2R_{\theta}(R_{\theta}(9-7R_{\theta}(1-R_{\phi})-11R_{\phi})$
@@
$\tilde{A}_{(3,0)}^{\text{u}}=$ $\tilde{A}_{(3,0)}-R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})$
×$\times((R_{\theta}(1-R_{\phi})+R_{\phi})(2-R_{\theta}(1-R_{\phi})-R_{\phi})$
@@
$\tilde{A}_{(3,1)}^{\text{u}}=$ $\tilde{A}_{(3,1)}-6R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})(R_{\theta}(1-R_{\phi})+R_{\phi}$
@@
$\tilde{A}_{(3,2)}^{\text{u}}=$ $\tilde{A}_{(3,2)}-12R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})(R_{\theta}(1-R_{\phi})$
@@
$\tilde{A}_{(3,3)}^{\text{u}}=$ $\tilde{A}_{(3,3)}-2R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}(2\beta)\,,$
$\tilde{A}_{(0,2)}^{\text{u}}=$ $\tilde{A}_{(0,3)}^{\text{u}}=\tilde{A}_{(1,2)}^{\text{u}}=\tilde{A}_{(1,3)}^{\text{u}}=\tilde{A}_{(2,3)}^{\text{u}}=0\,.$

### v2

@@ C.1 Identical contacts
General formulas for the conductance and the noise intensity are obtained by averaging the expressions  (9), (10), (11) and (12) in the main text over the energy, using expressions for the amplitude (34). This result was obtained early in Ref. Niyazov et al. (2024) in different form. It is convenient to write the resulting formulas, using the parameters $\lambda,$ $R_{\theta}$ and $R_{\phi}$ instead of the parameters $t,$ $\theta$ and $\phi$, according to the following definitions:
@@
${\mathcal{T}}^{\mathrm{HI}}=C\sum_{m,n=0}^{1}A_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,$, (59)
${\mathcal{T}}_{2}^{\mathrm{HI}}=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,$, (60)
@@
[…]
@@ C.2 Non-identical contacts
$t_{\mathrm{R,L}}$ $=te^{\pm\beta}\,,\quad|\beta|\leq\lambda$ (63)
$t_{\mathrm{R}}t_{\mathrm{L}}$ $=e^{-2\lambda}\,$,
$t_{\mathrm{R}}/t_{\mathrm{L}}$ $=e^{2\beta}\,.$
In this case we may still use the expressions (59), (41), but the coefficients should be redefined, $A_{(m,n)}\to A_{(m,n)}^{\mathrm{u}}$, $\tilde{A}_{(m,n)}\to\tilde{A}_{(m,n)}^{\mathrm{u}}$, as follows
$C^{\mathrm{u}}=$ $C\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)\,$, (64)
$A_{(0,0)}^{\mathrm{u}}=$ $A_{(0,0)}\,$,
$A_{(0,1)}^{\mathrm{u}}=$ $0\,$,
$A_{(1,0)}^{\mathrm{u}}=$ $A_{(1,0)}+R_{\theta}\sinh^{2}\beta\,$,
$A_{(1,1)}^{\mathrm{u}}=$ $A_{(1,1)}+2R_{\theta}\sinh^{2}\beta\,$,
$\tilde{C}^{\mathrm{u}}=$ $\tilde{C}\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)^{2}\,$,
$\tilde{A}_{(0,0)}^{\mathrm{u}}=$ $\tilde{A}_{(0,0)}+2R_{\theta}(2-R_{\theta})\sinh^{2}\beta\,$,
$\tilde{A}_{(0,1)}^{\mathrm{u}}=$ $\tilde{A}_{(0,1)}\,$,
[…]
@@
$\tilde{A}_{(1,1)}^{\mathrm{u}}=$ $\tilde{A}_{(1,1)}+R_{\theta}^{2}\sinh^{2}(2\beta)\,$,
$\tilde{A}_{(2,0)}^{\mathrm{u}}=$ $\tilde{A}_{(2,0)}+R_{\theta}(R_{\theta}(1-R_{\phi})+R_{\phi})$
[…]
@@
$\tilde{A}_{(2,1)}^{\mathrm{u}}=$ $\tilde{A}_{(2,1)}-4R_{\theta}(2-R_{\theta})(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}\beta$
[…]
@@
$\tilde{A}_{(3,0)}^{\mathrm{u}}=$ $\tilde{A}_{(3,0)}-R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})$
$\times((R_{\theta}(1-R_{\phi})+R_{\phi})(2-R_{\theta}(1-R_{\phi})-R_{\phi})$
@@
$\tilde{A}_{(3,1)}^{\mathrm{u}}=$ $\tilde{A}_{(3,1)}-6R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})(R_{\theta}(1-R_{\phi})+R_{\phi}$
@@
[…]
@@
$\tilde{A}_{(3,3)}^{\mathrm{u}}=$ $\tilde{A}_{(3,3)}-2R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}(2\beta)\,$,
$\tilde{A}_{(0,2)}^{\mathrm{u}}=$ $\tilde{A}_{(0,3)}^{\mathrm{u}}=\tilde{A}_{(1,2)}^{\mathrm{u}}=\tilde{A}_{(1,3)}^{\mathrm{u}}=\tilde{A}_{(2,3)}^{\mathrm{u}}=0\,.$